### v2

# Integrating Deep-Learning-Based Magnetic Model and Non-Collinear Spin-Constrained Method: Methodology, Implementation and Application

Abstract
We propose a non-collinear spin-constrained method that generates training data for deep-learning-based magnetic model, which provides a powerful tool for studying complex magnetic phenomena that requires large-scale simulations at the atomic level. First, we propose a projection method for calculating atomic magnetic moments by applying a radial truncation to numerical atomic orbitals. We then implement a Lagrange multiplier method that yields the magnetic torques of atoms by constraining the magnitude and direction of atomic magnetic moments. The method is implemented in ABACUS with both plane wave basis and numerical atomic orbital basis. We benchmark the iron (Fe) systems and analyze differences from calculations with the plane wave basis and numerical atomic orbitals basis in describing magnetic energy barriers. Based on more than 30,000 first-principles data with the information of magnetic torque, we train a deep-learning-based magnetic model DeePSPIN for the Fe system. By utilizing the model in large-scale molecular dynamics simulations, we successfully predict Curie temperatures of $\alpha$-Fe close to experimental values.
I Introduction
The study of magnetic materials has been a cornerstone of condensed matter physics, with implications ranging from fundamental science to technological applications. Density Functional Theory (DFT) Hohenberg and Kohn (1964); Kohn and Sham (1965) has emerged as a powerful tool for understanding the electronic and magnetic properties of these materials Szilva et al. (2023). However, accurate description of excited magnetic states within DFT has remained a formidable challenge due to the complex interplay between electron and lattice behaviors.
The development of constrained Density Functional Theory (cDFT) has marked a significant advancement in this field  Dederichs et al. (1984a); Kaduk et al. (2012). In the pioneering work of Dederichs et al. in 1984 Dederichs et al. (1984a), DFT was extended to arbitrary constraints through the introduction of the Lagrange multiplier method. Since then, several cDFT methods have been used to study excited states of charge distribution Dederichs et al. (1984a); Wu and Van Voorhis (2005) or magnetization Stocks et al. (1998); Kurz et al. (2004); Cuadrado et al. (2018); Gebauer and Baroni (2000); Ma and Dudarev (2015); Hegde et al. (2020a); Cai et al. (2023). Introducing penalty functions is commonly adopted to implement any constraint. Gebauer and Baroni (2000); Ma and Dudarev (2015). Wu and Van Voorhis (2005) Wu and Van Voorhis (2005) introduced an efficient algorithm, which allows for the use of a double-loop method to find the effective Lagrange multiplier that satisfies the constraints. In addition, the cDFT methods have been implemented based on different basis sets, such as the real-space grid Gebauer and Baroni (2000), the full-potential linearized augmented plane-wave (FLAPW) Kurz et al. (2004), the projector augmented-wave (PAW) Ma and Dudarev (2015); Hegde et al. (2020a); Cai et al. (2023), and the numerical atomic orbitals (NAOs) Cuadrado et al. (2018). These efforts provide powerful tools in understanding charge and magnetization fluctuations in solids, predicting spin-dependent phenomena, and characterizing electron transfer reactions in molecules Kaduk et al. (2012); Stocks et al. (1998); Kurz et al. (2004); Wu and Van Voorhis (2005); Ma and Dudarev (2015); Cuadrado et al. (2018); Hegde et al. (2020a); Cai et al. (2023).
Reviewing the numerous spin-cDFT methods, one of main differences lies in the definitions of atomic magnetic moments. One category is the Mulliken population method Cuadrado et al. (2018), which, while maintaining the sum rule, may introduce contributions from non-local orbitals. Another approach is based on the local orbital projection Hegde et al. (2020a); Cai et al. (2023), which preserves better the locality of magnetic moments. In this work, we implement a spin-constrained method with both plane-wave (PW) basis and numerical atomic orbital basis within the DFT package ABACUS Chen et al. (2010); Li et al. (2016); Lin et al. (2020); Zhou et al. (2025). Benefiting from the locality of NAOs, magnetic moments are obtained through the projection of atomic orbitals. By appropriately truncating the local atomic orbitals and smoothing near the truncation, we find that the modulated atomic orbitals provide satisfactory performance in stability and convergence during cDFT calculations. The spin-cDFT is fully optimized using the double-loop approach Wu and Van Voorhis (2005). Importantly, the optimized Lagrange multipliers can serve as magnetic torques for atoms.
In recent years, the integration of deep learning models with DFT has offered unprecedented opportunities for transferring first-principles accuracy to larger scales Zhang et al. (2024, 2023); Qi et al. (2024); Chen and Ong (2022); Deng et al. (2023); Batatia et al. (2023); Choudhary et al. (2023); Choudhary and DeCost (2021). The available databases Jain et al. (2013); Talirz et al. (2020); ato ; Barroso-Luque et al. (2024) contain a large amount of DFT data that can be used to train machine learning models for different purposes. Liu et al. (2024); Wu et al. (2023a, b). Recently, some magnetic models Rinaldi et al. (2024); Yang et al. (2024) have emerged that incorporate the degrees of freedom of spin into the existing neural networks and can perform large-scale magnetic dynamics simulations. However, these models suffer from a lack of data since most existing DFT data in datasets are non-magnetic or only include ground-state magnetic results. Ground-state data cannot meet the sampling requirements of magnetic models with an additional degree of freedom spin. From this perspective, spin-cDFT methods not only act as powerful first-principles tools for studying magnetic excited states at the atomic scale but also provide training data for machine-learning-based magnetic models. Additionally, incorporating magnetic torque is advantageous for magnetic models that can describe magnetic field interactions. This requires that the analytical form of the magnetic torque can be satisfied through the definition of the constraints. However, most existing methods of cDFT, especially those with penalty functions, do not meet this requirement. In this work, we implement a spin-cDFT method within the ABACUS package, which provides energies, atomic forces, stresses, and magnetic torques for any magnetic excited state, making it a suitable data engine for magnetic models. As an example, we train a DeePSPIN model for the iron system and predict the Curie temperature through molecular dynamics simulations.
The content of this paper is as follows. Sec. II details the theoretical foundations, including non-collinear spin, projection methods, and spin-constrained DFT. Sec. III.2 discusses the modulated NAOs, as well as the behavior of atomic magnetic moments based on orbital projections. Sec. III.3 verifies the correctness of the implementation through finite difference tests. Sec. III.4 introduces the properties of various magnetic excited states of pure Fe calculated using spin-cDFT. Sec. III.5 shows how to train a DeePSPIN model starting from the spin-cDFT data of Fe, and provides an accurate Curie temperature through molecular dynamic simulation. Finally, Sec. IV summarizes the work.
II Method
II.1 Non-Collinear Spin
For systems with non-collinear spin configurations, where the spin is not uniformly aligned along a single axis, the electronic wavefunction is expressed as a spinor $\psi=\{\psi^{\uparrow},\psi^{\downarrow}\}$. In the context of non-collinear magnetism within density functional theory, the Kohn-Sham equations for a two-component spinor wave function $\psi$ can be written as Kohn and Sham (1965)
$\left[-\frac{\hbar^{2}}{2m}\nabla^{2}+V_{\text{eff}}(\mathbf{r})\right]\psi_{i}(\mathbf{r})=\epsilon_{i}\psi_{i}(\mathbf{r}),$ (1)
where $V_{\text{eff}}$ is the effective potential, including the Hartree term $V_{\text{H}}$, the external potential term $V_{\text{ext}}$ and the exchange-correlation (XC) term $V_{\text{xc}}$. In addition, the generalized charge density can be defined by introducing two-component spin space and the formula is
$\bar{\rho}=\frac{1}{2}(\rho+\sum_{p}m_{p}\cdot\bm{\sigma}^{p})=\frac{1}{2}\left(\begin{array}[]{cc}\rho+m_{z}&m_{x}-im_{y}\\ m_{x}+im_{y}&\rho-m_{z}\end{array}\right),$ (2)
where
$\rho(\mathbf{r})=\sum_{i}f_{i}\psi_{i}^{\dagger}(\mathbf{r})\psi_{i}(\mathbf{r})$ (3)
is charge density and the spin density vector $\mathbf{m}=(m_{x},m_{y},m_{z})$ is
$m_{p}(\mathbf{r})=\sum_{i}f_{i}\psi_{i}^{\dagger}(\mathbf{r})\bm{\sigma}^{p}\psi_{i}(\mathbf{r}).$ (4)
Here $\psi_{i}$ is two component spinor wavefunction for $i$-th band with occupation number $f_{i}$. $p$ takes values {1, 2, 3}, representing the components of spin density $\mathbf{m}$ along the $x$, $y$ and $z$ directions. At the same time, it also denotes the $p$-th Pauli matrix $\bm{\sigma}^{p}$.
The magnetic effects in the Hamiltonian can be categorized into two parts. First, the electron-electron interaction, which is encompassed within the exchange-correlation term, directly influences the self-consistent solution process and necessitates the careful selection of appropriate functionals. Second, the spin-orbit coupling (SOC) effect, arising from relativistic effects, is introduced through fully relativistic pseudopotentials and is computed within the non-local components of the pseudopotential.
In the computation of the XC functional, the treatment of non-collinear spins can be simplified by using the local density approximation (LDA), which reduces the problem to calculating collinear spins at each grid point. For local reference systems used to calculate the XC functional, the local charge density can be transformed into the diagonal matrix form of the generalized charge density as
$\bar{\rho}^{\prime}=\left(\begin{array}[]{cc}\rho_{+}&0\\ 0&\rho_{-}\end{array}\right),$ (5)
where $\rho_{+}$ and $\rho_{-}$ are defined as
$\rho_{\pm}=\frac{1}{2}(\rho\pm|\mathbf{m}|)=\frac{1}{2}\left(\rho\pm\sqrt{m_{x}^{2}+m_{y}^{2}+m_{z}^{2}}\right).$ (6)
Taking into account the non-collinear electron spins within the framework of the aforementioned local spin density approximation (LSDA) Miehlich et al. (1989), the expressions for the exchange-correlation energy functional takes the form of
[…]
and the corresponding potential is
$V_{\mathrm{XC}}(\mathbf{r})$ $=\frac{\delta E_{\mathrm{XC}}}{\delta\rho(\mathbf{r})}$ (8)
$=\varepsilon_{\mathrm{XC}}(\rho(\mathbf{r}),|\mathbf{m}(\mathbf{r})|)+\rho(\mathbf{r})\left[\frac{|\partial\varepsilon_{\mathrm{XC}}(\rho(\mathbf{r}),|\mathbf{m}(\mathbf{r})|)}{\partial\rho(\mathbf{r})}\right].$
The computation of the XC magnetic field $\mathbf{b}(\mathbf{r})$ can be achieved by employing the chain rule to differentiate the modulus of the spin density, ensuring that $\mathbf{b}(\mathbf{r})$ is collinear with the magnetization vector $\mathbf{m}(\mathbf{r})$. The formula is
$\mathbf{b}(\mathbf{r})$ $=-\frac{\delta E_{\mathrm{XC}}}{\delta\mathbf{m}(\mathbf{r})}=-\frac{\delta|\mathbf{m}(\mathbf{r})|}{\delta\mathbf{m}(\mathbf{r})}\frac{\delta E_{\mathrm{XC}}}{\delta|\mathbf{m}(\mathbf{r})|}$ (9)
$=-\widehat{\mathbf{m}}(\mathbf{r})\rho(\mathbf{r})\left[\frac{\partial\varepsilon_{\mathrm{XC}}(\rho(\mathbf{r}),|\mathbf{m}(\mathbf{r})|)}{\partial|\mathbf{m}(\mathbf{r})|}\right].$
[…]
The spin-orbit coupling and relativistic effects are incorporated using norm-conserving pseudopotentials Hamann et al. (1979); Troullier and Martins (1991) within the Kleinman-Bylander (KB) form  Hemstreet et al. (1993); Theurich and Hill (2001); Kleinman (1980); Bachelet and Schlüter (1982). Additionally, spin-orbit terms are directly included in both plane-wave basis sets Corso and Conte (2005) and numerical atomic orbitals basis sets Cuadrado and Cerdá (2012) for self-consistent field calculations.
II.2 Projection Methods
As the core physical quantity in spin-cDFT, the atomic magnetic moment is crucial for accurate algorithmic implementation. Typically, one can define atomic magnetic moments in three ways: by partitioning the spin density Bader and Molecules (1990); Fonseca Guerra et al. (2004); Hirshfeld (1977), through wavefunction analysis Mulliken (1955); Löwdin (1970), and via subspace projection Hegde et al. (2020b).
A straight definition belonging to the first kind is the “spherical definition method”, in which the spin density is integrated over a spherical volume around the nucleus that is defined by the window function $\Theta(\mathbf{r};\bm{\tau}_{I})$. For example
$\mathbf{M}_{I}=\int\mathbf{M}(\mathbf{r})\Theta(r^{\text{cut}}-|\mathbf{r}-\bm{\tau}_{I}|)\mathrm{d}\mathbf{r},$ (10)
where $\bm{\tau}_{I}$ denotes the position of the nucleus $I$ and $r^{\mathrm{cut}}$ is the cutoff radius, $\Theta$ is the Heaviside step function or other smoothed window function.
More sophisticated strategy requires the conservation on quantities to be partitioned,
$\mathbf{M}_{I}\left(\mathbf{r}\right)=\mathbf{M}\left(\mathbf{r}\right)w_{I}\left(\mathbf{r}\right),$ (11)
in which
$w_{I}\left(\mathbf{r}\right)=\frac{P_{I}\left(\mathbf{r}\right)}{\sum_{J}{P_{J}\left(\mathbf{r}\right)}}.$ (12)
$w_{I}(\mathbf{r})$ is the weighting function and partition function $P_{I}(\mathbf{r})$ approaches to 1 near $I$-th atom and 0 otherwise. A famous example of a more sophisticated partitioning strategy is called Atom-In-Molecule (AIM), or well-known Bader charge analysis Bader and Molecules (1990), which defines atomic regions by zero-flux surfaces in the gradient of the electron density, Voronoi tessellation Fonseca Guerra et al. (2004) partitions space into polyhedral cells around each atom based on proximity. The Bader charge analysis provides a mathematically rigorous definition of atomic boundaries based on electron density, yielding charges that are invariant to the choice of basis set and more reflective of the true electronic structure. However, this method would fail in cases where there are zero or more than one atoms appear in one Voronoi cell, and lack of decomposition of the orbits.
Mulliken Mulliken (1955) and Löwdin Löwdin (1970) charge analysis methods are well-established for partitioning electron density within a molecule to assign atomic charges. The formula for calculating the atomic magnetization from Mulliken charge of $I$-th atom is given by
$\mathbf{M}_{I}^{\mathrm{Mulliken}}=\sum_{\mu\in I}\sum_{\mathbf{k}}\sum_{\nu}\bm{\sigma}\cdot\bm{\rho}_{\mu\nu}(\mathbf{k})S_{\mu\nu}(\mathbf{k}).$ (13)
Similarly, the formula for the Löwdin charge is:
$\mathbf{M}_{I}^{\mathrm{L\ddot{o}wdin}}=\sum_{\mu\in I}\sum_{\mathbf{k}}\sum_{\nu}\bm{\sigma}\cdot S_{\mu\nu}^{1/2}(\mathbf{k})\bm{\rho}_{\mu\nu}(\mathbf{k})S_{\mu\nu}^{1/2}(\mathbf{k}),$ (14)
where $\bm{\sigma}$ is the Pauli matrix to decompose reciprocal space density matrix $\bm{\rho}_{\mu\nu}(\mathbf{k})$, and $S_{\mu\nu}(\mathbf{k})$ is the overlap integral matrix between basis functions $\mu$ and $\nu$ on this atom. These two charge analysis methods are powerful tools for approximating electron distribution by assigning shared electrons between atoms based on orbital overlaps, thereby offering a rudimentary yet quick insight into molecular charge states. However, their reliance on the choice of basis sets and the arbitrary nature of electron partitioning, which overlooks electronegativity differences, often results in charges that lack physical accuracy and can vary significantly with computational parameters.
The subspace projection method requires a rational construction of projection operators in which the projection function should be of physical meaning. For example in projected augmented wave (PAW) formulism Blöchl (1994), the PAW projector $\{|p_{I\mu}\rangle\}$ which holds the orthonormality with the pseudo partial wave that is connected with the all-electron one $\{|\phi_{\mu}\rangle\}$ is a natural choice Ma and Dudarev (2015); Hegde et al. (2020a). Atomic magnetic moments is defined as
$\mathbf{M}^{p}_{I}=\sum_{\sigma\sigma^{\prime}}\sum_{\mu\nu}\bm{\sigma}^{p\sigma\sigma^{\prime}}D_{I\mu\nu}^{\sigma\sigma^{\prime}}\Omega_{\mu\nu},$ (15)
in which the scalar $\sigma$ and $\sigma^{\prime}$ refer to the spin index (up and down) and the $\bm{\sigma}^{p}$ is the $p$-th Pauli matrix. $D_{I\mu\nu}^{\sigma\sigma^{\prime}}$ is the one-center density matrix representation of pseudo partial wave centered at the $I$-th atom,
$D_{I\mu\nu}^{\sigma\sigma^{\prime}}=\sum_{n}f_{n}^{\sigma}\langle\psi_{n}^{\sigma}|p_{I\mu}\rangle\langle p_{I\nu}|\psi_{n}^{\sigma^{\prime}}\rangle,$ (16)
$\Omega_{\mu\nu}$ is the all-electron partial wave matrix elements of the cutoff-sphere integral
$\Omega_{\mu\nu}=\int d^{3}\mathbf{r}\phi_{\mu}(\mathbf{r})\phi_{\nu}(\mathbf{r})\Theta(r^{\mathrm{cut}}-|\mathbf{r}-\mathbf{\bm{\tau}}_{I}|).$ (17)
This algorithm effectively utilizes the PAW projection orbitals’ ability to accurately describe different electronic angular momentum orbitals, while also leveraging spherical truncation to preserve locality, making it a highly efficient method for defining atomic magnetic moments.
In this work, we propose a projection method under pseudopotential formalism for calculating atomic magnetization, we employ valence electron orbital local density projection operators
$\hat{P}_{Ilmm^{\prime}}=|\alpha_{Ilm}\rangle\langle\alpha_{Ilm^{\prime}}|$ (18)
to estimate atomic magnetic moments $\mathbf{M}_{I}$, where $\alpha_{Ilm}$ is the projection orbital distinguished by the angular momentum quantum number $l$ and the magnetic quantum number $m$. Within this framework, the sum of occupancy of the valence electrons $\sum_{m}n^{\sigma\sigma^{\prime}}_{Ilm}$ with identical angular momentum $l$ and spin index $\sigma,\sigma^{\prime}$ corresponds to the trace of the projected density matrix $\text{Tr}(n^{\sigma\sigma^{\prime}}_{Ilmm^{\prime}})$, where
$n^{\sigma\sigma^{\prime}}_{Ilmm^{\prime}}=\sum_{i}f_{i}\langle\psi^{\sigma}_{i}|\hat{P}_{Ilmm^{\prime}}|\psi^{\sigma^{\prime}}_{i}\rangle.$ (19)
The atomic magnetic moment along the direction $p$ is calculated by summing the diagonal elements of the occupation matrix for each angular momentum
$\mathbf{M}_{I}^{p}=\sum_{\sigma\sigma^{\prime}}{{\sum_{lmm^{\prime}}{\bm{\sigma}_{\sigma\sigma^{\prime}}^{p}n_{Ilmm^{\prime}}^{\sigma\sigma^{\prime}}\delta_{mm^{\prime}}}}},$ (20)
where $p$ refers to $x,y,z$ components.
The accuracy and locality of the projection depend upon the orbitals $|\alpha_{Ilm}\rangle$ used to construct the projection operators $\hat{P}_{Ilmm^{\prime}}$, which are composed of radial distribution functions and spherical harmonics
$\alpha_{Ilm}(\mathbf{r})=\alpha_{l}(|\mathbf{r}-\bm{\tau}_{I}|)Y_{lm},$ (21)
where the radial functions $\alpha(r)$ are constructed to meet three criteria: (1) maximized efficiency for extracting valence electron information of interest, (2) normalization, and (3) smoothness at the boundary to avoid numerical error. Practically, we construct smooth modulation orbitals (SMOs) by truncating those $\zeta$ functions of ABACUS numerical atomic orbitals (NAOs) Lin et al. (2021); Chen et al. (2010) that collect majority of valence electron information over various reference systems, then smoothed by a function centered at the boundary
$g(r;\sigma)=\left\{\begin{matrix}1-\exp\left(-\frac{(r-r_{m})^{2}}{2\sigma^{2}}\right)&r<r_{m}\\ 0&r\geq{}r_{m}\end{matrix}\right.,$ (22)
where $r$ is the distance to the atom, $r_{m}$ is the artificially chosen modulation radius. Spreading parameter $\sigma$ is solved in an iterative way, that can maximize the overlap integral between the SMOs and NAOs under the constraint of normalization of the SMOs themselves
$\alpha(r)=\left.\frac{\chi(r)g(r;\sigma)}{\langle\chi(r)g(r;\sigma)|\chi(r)g(r;\sigma)\rangle^{1/2}}\right|_{\frac{\partial\langle\alpha|\chi\rangle}{\partial\sigma}=0},$ (23)
where $\chi(r)$ are the radial parts of $\zeta$ functions in NAOs. Using SMOs as projection orbitals allows for the representation of charge distributions across different angular momentum orbitals within an atom’s localized environment. Since the projection orbitals are designed to be strongly localized, inter-atomic overlap can be neglected. Furthermore, the orthogonality of projection orbitals with different angular momenta is preserved during modulation, resulting in a simplified form for the projection operator.
The onsite occupation matrix, which quantifies the localized atomic charge occupancy, can be obtained through the projection of wave functions onto SMOs. In NAO formulation for periodic systems, the onsite occupation matrix can be written as
$n^{\sigma\sigma^{\prime}}_{Ilmm^{\prime}}$ $=\sum_{n\mathbf{k}}f_{n\mathbf{k}}\langle\psi_{n\mathbf{k}}^{\sigma}|\alpha_{Ilm}\rangle\langle\alpha_{Ilm^{\prime}}|\psi_{n\mathbf{k}}^{\sigma^{\prime}}\rangle$ (24)
$=\sum_{\mathbf{R}}\sum_{\mu\nu}\rho_{\mu\nu}^{\sigma\sigma^{\prime}}(\mathbf{R})\langle\phi^{\mathbf{0}}_{\mu}|\alpha_{Ilm}\rangle\langle\alpha_{Ilm^{\prime}}|\phi^{\mathbf{R}}_{\nu}\rangle,$
where
$|\psi_{n\mathbf{k}}^{\sigma}\rangle=\sum_{\mu}c^{\sigma}_{n\mu}(\mathbf{k})\phi_{\mu}(\mathbf{k})$ (25)
represents the eigenvector of $n$-th band at the first Brillouin zone sampling point $\mathbf{k}$, and the spin index is $\sigma$. The parameter $f_{n\mathbf{k}}$ depicts occupation numbers of electrons. $\phi^{\mathbf{0}}_{\mu}=\phi_{\mu}(\mathbf{r}-\bm{\tau}_{\mu})$ and $\phi^{\mathbf{R}}_{\nu}=\phi_{\nu}(\mathbf{r}-\mathbf{R}-\bm{\tau}_{\nu})$ are numerical orbital functions. $\rho_{\mu\nu}^{\sigma\sigma^{\prime}}(\mathbf{R})$ is the real space density matrix that takes the form of
$\rho_{\mu\nu}^{\sigma\sigma^{\prime}}(\mathbf{R})=\frac{1}{N_{\mathbf{k}}}\sum_{n\mathbf{k}}f_{n\mathbf{k}}c_{n\mu}^{\sigma*}(\mathbf{k})c_{n\nu}^{\sigma^{\prime}}(\mathbf{k})e^{i\mathbf{k\cdot R}}.$ (26)
For plane wave basis, the onsite occupation matrix $n^{\sigma\sigma^{\prime}}_{Ilmm^{\prime}}$ can be written as
$n^{\sigma\sigma^{\prime}}_{Ilmm^{\prime}}$ $=\sum_{n\mathbf{k}}f_{n\mathbf{k}}S_{In\mathbf{k}lm}^{\sigma*}S_{In\mathbf{k}lm^{\prime}}^{\sigma^{\prime}},$ (27)
where
$S_{In\mathbf{k}lm}^{\sigma}=\sum_{\mathbf{G}}{\alpha_{Ilm}^{*}(\mathbf{G})c_{n\mathbf{k}}^{\sigma}(\mathbf{G})}$ (28)
is the plane wave basis representation of the overlap between $|\psi_{n\mathbf{k}}^{\sigma}\rangle$ and SMOs, $c_{n\mathbf{k}}^{\sigma}(\mathbf{G})$ and $\alpha_{Ilm}(\mathbf{G})$ are coefficients of plane wave expansion of $|\psi_{n\mathbf{k}}^{\sigma}\rangle$ and SMOs, respectively. One can refer to Ref. King-Smith et al., 1991 for the details of SMOs expansion.
II.3 Spin-Constrained DFT
Figure 1: Framework of variant Lagrange quantities spin-cDFT algorithm implemented in ABACUS for both PW and NAOs basis. The electron density is $\rho$ and the spin density vector is $\mathbf{m}$. The atomic magnetization is $\mathbf{M}$.
Utilizing the established Lagrange formalism Dederichs et al. (1984b); Cai et al. (2023), the challenge pertaining to the restriction of atomic magnetic moments is reformulated as an endeavor to ascertain the equilibrium points of the pertinent function. To this end, a Lagrange multiplier is introduced into the energy functional of the system:
$E_{c}=E_{\text{KS}}$ $\big{(}\rho(\mathbf{r}),\mathbf{m}(\mathbf{r})\big{)}$ (29)
$-\sum_{I}\bm{\lambda}_{I}\cdot\Big{(}\mathbf{M}_{I}\big{(}\mathbf{m}(\mathbf{r})\big{)}-\mathbf{M}_{I,\text{target}}\Big{)},$
where $E_{\text{KS}}$ is the Kohn-Sham energy with charge density $\rho(\mathbf{r})$ and spin density $\mathbf{m}(\mathbf{r})$, $\bm{\lambda}_{I}$ is the Lagrange multiplier, which can be treated as the magnetic torque under this constraint. $\mathbf{M}_{I}$ and $\mathbf{M}_{I,\text{target}}$ are the atomic spin moment and the target atomic spin moment for atom $I$, respectively. The stationary point problem within the Lagrange formalism can be solved iteratively by minimizing $E_{c}$ with respect to $\bm{\lambda}_{I}$, $\rho(\mathbf{r})$ and $\mathbf{m}(\mathbf{r})$ Wu and Van Voorhis (2005).
II.3.1 Numerical Atomic Orbital Basis
The penalty Hamiltonian term in NAOs basis in real space can be derived from the Lagrange function as
$H_{\mu\nu}^{\bm{\lambda},\sigma\sigma^{\prime}}(\mathbf{R})$ $=\frac{\partial E_{c}(\rho,\{\bm{\lambda}_{I}\},\{\mathbf{M}_{I}\})}{\partial\rho^{\sigma\sigma^{\prime}}_{\mu\nu}(\mathbf{R})}$ (30)
$=\sum_{I}\frac{\partial E_{c}(\rho,\bm{\lambda}_{I},\mathbf{M}_{I})}{\partial\mathbf{M}_{I}}\frac{\partial\mathbf{M}_{I}}{\partial\rho^{\sigma\sigma^{\prime}}_{\mu\nu}(\mathbf{R})}$
$=\sum_{I}\sum_{p}\frac{\partial({\lambda}_{I}^{p}M_{I}^{p})}{\partial\rho^{\sigma\sigma^{\prime}}_{\mu\nu}(\mathbf{R})}$
$=\sum_{I}\sum_{p}{\lambda}_{I}^{p}\frac{\partial M_{I}^{p}}{\partial\rho^{p}_{\mu\nu}(\mathbf{R})}\frac{\partial\rho^{p}_{\mu\nu}(\mathbf{R})}{\partial\rho^{\sigma\sigma^{\prime}}_{\mu\nu}(\mathbf{R})}.$
Here
$\rho^{p}_{\mu\nu}(\mathbf{R})=\sum_{\sigma\sigma^{\prime}}\bm{\sigma}^{p}\rho^{\sigma\sigma^{\prime}}_{\mu\nu}(\mathbf{R}),$ (31)
where $\mathbf{R}$ is the lattice vector between basis functions $\mu$ and $\nu$. The last term in the chain derivative on the right-hand side is Pauli matrix, the rest of the term can continue to expand the calculation by Eq. 20 as
$\sum_{lm}\frac{\partial M_{I}^{p}}{\partial n^{p}_{Ilm}}\frac{\partial n^{p}_{Ilm}}{\partial\rho^{p}_{\mu\nu}(\mathbf{R})}=\sum_{lm}\langle\phi^{\mathbf{0}}_{\mu}|\alpha_{Ilm}\rangle\langle\alpha_{Ilm}|\phi^{\mathbf{R}}_{\nu}\rangle.$ (32)
The penalty term of Hamiltonian has the form of
$H_{\mu\nu}^{\bm{\lambda},\sigma\sigma^{\prime}}(\mathbf{R})=\sum_{I}f(I,\sigma\sigma^{\prime})\sum_{lm}\langle\phi^{\mathbf{0}}_{\mu}|\alpha_{Ilm}\rangle\langle\alpha_{Ilm}|\phi^{\mathbf{R}}_{\nu}\rangle,$ (33)
where the 2$\times$2 parameter matrix $f(I,\sigma\sigma^{\prime})$ from Pauli matrix is
$f(I,\sigma\sigma^{\prime})=\left(\begin{array}[]{cc}\lambda_{I}^{x}&\lambda_{I}^{x}+i\lambda_{I}^{y}\\ \lambda_{I}^{x}-i\lambda_{I}^{y}&-\lambda_{I}^{z}\end{array}\right).$ (34)
This modified Hamiltonian ensures that the results satisfy the constraints and contribute accordingly to the total energy, and the contribution of energy from the penalty term as follows:
$E^{\bm{\lambda}}=\sum_{i}f_{i}\langle\psi_{i}|\hat{h}^{\bm{\lambda}}|\psi_{i}\rangle-\sum_{I}\bm{\lambda}_{I}\cdot\mathbf{M}_{I,target},$ (35)
where $\hat{h}^{\bm{\lambda}}$ is symbol of penalty operator.
The correction terms of atomic force and lattice stress can be calculated by the derivative of the energy term in Eq. 35. The expression of penalty force in the direction along $p=x,y,z$ is
$\mathbf{F}^{\bm{\lambda},p}_{I}=-2\sum_{\mathbf{R}}$ $\sum_{\mu\nu}\sum_{\sigma\sigma^{\prime}}\rho_{\mu\nu}^{\sigma\sigma^{\prime}}(\mathbf{R})\times$ (36)
$\sum_{\mathbf{R}^{\prime}}\sum_{lm}\langle\phi_{\mu}^{\mathbf{0}}|\alpha_{Ilm}^{\mathbf{R}^{\prime}}\rangle f(I,\sigma\sigma^{\prime})\frac{\partial}{\partial\bm{\tau}^{p}_{I}}\langle\alpha_{Ilm}^{\mathbf{R}^{\prime}}|\phi_{\nu}^{\mathbf{R}}\rangle,$
where $\tau^{p}_{I}$ is the $p$-component coordinates of $I$-th atom $\bm{\tau}$. The penalty stress term is
$\Sigma^{\bm{\lambda}}_{\alpha\beta}=-\frac{1}{\Omega}$ $\sum_{\mathbf{R}}\sum_{\mu\nu}\sum_{\sigma\sigma^{\prime}}\rho_{\mu\nu}^{\sigma\sigma^{\prime}}(\mathbf{R})\sum_{I}\sum_{\mathbf{R}^{\prime}}\sum_{lm}$ (37)
$\Big{(}\frac{\partial}{\partial\tau^{\alpha}_{\mu I}}\langle\phi_{\mu}^{\mathbf{0}}|\alpha_{Ilm}^{\mathbf{R}^{\prime}}\rangle\tau^{\beta}_{\mu I}f(I,\sigma\sigma^{\prime})\langle\alpha_{Ilm}^{\mathbf{R}^{\prime}}|\phi_{\nu}^{\mathbf{R}}\rangle$
$+\langle\phi_{\mu}^{\mathbf{0}}|\alpha_{Ilm}^{\mathbf{R}^{\prime}}\rangle f(I,\sigma\sigma^{\prime})\frac{\partial}{\partial\tau^{\alpha}_{I\nu}}\langle\alpha_{Ilm}^{\mathbf{R}^{\prime}}|\phi_{\nu}^{\mathbf{R}}\rangle\tau^{\beta}_{I\nu}\Big{)}$
where $\alpha,\beta$ are indices of stress $3\times 3$ tensor.
II.3.2 Plane Wave Basis
Given a large number of plane wave basis functions, directly constructing the full Hamiltonian matrix is generally impractical. Instead, iterative diagonalization methods are employed to solve the Kohn-Sham equations, which require only the computation of the action of the Hamiltonian operator on the wave function. As a result, within the plane-wave basis set, the formula for the Hamiltonian correction term at a specified $\mathbf{k}$-point is as follows
$\hat{h}^{\bm{\lambda}}|c^{\sigma}_{n\mathbf{k}}(\mathbf{G})\rangle=\sum_{I}\sum_{\sigma^{\prime}}f(I,\sigma\sigma^{\prime})\sum_{lm}S_{In\mathbf{k}lm}^{\sigma^{\prime}}|\alpha_{Ilm}(\mathbf{k+G})\rangle.$ (38)
where
$S_{In\mathbf{k}lm}^{\sigma}=\langle\alpha_{Ilm}(\mathbf{k+G})|c^{\sigma}_{n\mathbf{k}}(\mathbf{G})\rangle$ (39)
is the overlap of projected orbital and wave function.
The terms for the correction of atomic force and lattice stress can be also derived from the energy term of the Eq. 35 through its differentiation. The expression for atomic forces is
$\bm{F}^{\bm{\lambda},p}_{I}=-2\sum_{\mathbf{k}}$ $\sum_{I^{\prime}}\sum_{lm}\sum_{\sigma\sigma^{\prime}}f(I^{\prime},\sigma\sigma^{\prime})S_{I^{\prime}n\mathbf{k}lm}^{\sigma}$ (40)
$\langle\alpha_{I^{\prime}lm}(\mathbf{G+k})|(-j)(\mathbf{G+k})_{p}|c^{\sigma^{\prime}}_{n\mathbf{k}}(\mathbf{G})\rangle,$
and for stress is
$\Sigma^{\bm{\lambda}}_{\alpha\beta}=-\frac{2}{\Omega}\sum_{\mathbf{k}}\sum_{I^{\prime}}\sum_{lm}\sum_{\sigma\sigma^{\prime}}f(I^{\prime},\sigma\sigma^{\prime})S_{I^{\prime}n\mathbf{k}lm}^{\sigma}\frac{\partial S_{I^{\prime}n\mathbf{k}lm}^{\sigma^{\prime}}}{\partial\varepsilon_{\alpha\beta}},$ (41)
where $\Omega$ is the cell volume, $\varepsilon_{\alpha\beta}$ is the lattice strain tensor, and detail of last derivative term is similar with nonlocal pseudopotential term in Ref. Nielsen and Martin, 1985.
II.3.3 Workflow
We present a uniform flowchart of the specific implementation of spin-cDFT in Fig. 1 for both PW and NAO basis. As depicted in Fig. 1, this procedure necessitates an additional iteration within the self-consistent field (SCF) cycle of DFT to ascertain the fixed-point solution satisfying the spin constraints. The process involves solving the Kohn-Sham equations with a given set of Lagrange multipliers, which are then updated to minimize the deviation between the calculated and target magnetic moments. This minimization is typically achieved through a conjugate-gradient scheme or a similar optimization technique, ensuring that the magnetic moments converge to the targets within a predefined tolerance. The iterative loop continues, alternating between the solution of the Kohn-Sham equations and the optimization of the Lagrange multiplier, until the system reaches a self-consistent solution where the constraints are satisfied to a desired level of accuracy.
III Results and Discussions
III.1 Computational Details
Figure 2: Radial functions (RFs) of the first-$\zeta$ numerical atomic orbital of Fe before and after modulation, where (a) $p$ orbitals and (b) $d$ orbitals display the radial functions of the original numerical atomic orbitals with a cutoff radius $r_{c}$ of 6.0 Bohr, as well as the modulated orbital with the modification radius $r_{m}$ ranging from 1.0 to 5.0 Bohr. “p-1.0” represents the $p$ orbital modulated by Eq. 22 with $r_{m}=1.0$ Bohr.
All calculations in this work were performed with ABACUS, utilizing plane waves or numerical atomic orbitals Li et al. (2016) to expand wavefunctions. The pseudopotentials were sourced from the Pseudo-Dojo pseudopotential library Pseudo Dojo . The numerical atomic orbitals of ABACUS are generated by optimizing the spillage function averaged over structures; this strategy was originally proposed in the works of Chen et al. Chen et al. (2010) (denoted as v1.0) and improved by Lin et al. Lin et al. (2020) by including the kinetic terms (denoted as v2.0). In this work, a numerically further improved version is also used (denoted as v2.1). These v2.1 orbitals have not yet been officially released on the official website but the generation code is online available orb . This paper performs NAOs calculations using v2.0 and v2.1 orbitals, including Double-$\zeta$ plus polarization functions (DZP), and Triple-$\zeta$ plus double polarization functions (TZDP) with different cutoffs. Furthermore, for all calculations, the cutoff energy was set to 100 Ry, and the Brillouin zone was uniformly sampled at 0.14 Bohr-1 intervals. The projected orbitals used to calculate magnetic moments were modified numerical atomic orbitals, with a modulation radius of 3 Bohr. Each inner loop of $\bm{\lambda}$ optimization converges when the maximum atomic magnetism variation is less than $\delta\mathbf{M}<10^{-7}\mu_{B}$, while the SCF convergence until density difference $\delta\rho<10^{-6}$.
III.2 Projection Methods
As mentioned in Sec. II.2, we have developed an innovative algorithm designed to estimate and control atomic magnetic moments using localized orbital projection techniques. The key innovation of our method lies in the localized modulation of the numerical atomic orbital basis. We have benchmarked this algorithm on ferromagnetic (FM) and antiferromagnetic (AFM) iron (Fe) bulk.
Our first step involved comparing the modulated radial functions (RFs) of orbitals obtained through our algorithm with the original NAO RFs. It is important to distinguish the two properties discussed here: the cutoff radius $r_{c}$ and the modulation radius $r_{m}$ of the orbitals. The original NAO is zero beyond the cutoff radius, while the modulation radius refers to the range of the original NAO modulated by Eq. 22 for magnetic moment projection. In Fig. 2, we modify the shape of a NAO using Eq. 22 with $r_{m}$ ranging from 1 to 5 Bohr, where the $r_{c}$ of original NAO is 6.0 Bohr. As shown in Fig. 2(a), the modulated $p$ orbitals align closely with the original orbital shape if $r_{m}\geq 2$ Bohr, while the modulated $d$ orbitals remain consistent with the original shape for $r_{m}\geq 3$ Bohr. These modulated orbitals exhibit smoothness at $r_{m}$, with the first derivative approaching zero. For more localized $r_{m}$, the modulated orbitals retain the peak radius of the RFs, but their peak heights are significantly increased due to normalization constraints. Excessively large $r_{m}$ may introduce non-localized electron wavefunctions or portions of wavefunctions from localized orbitals of other atoms into the projection results, thereby preventing the projection from accurately representing atomic localized orbitals. On the other hand, excessively small $r_{m}$ can lead to significant distortion, as they cause a large deviation from the original atomic orbitals.
Next, we discuss the criteria for selecting appropriate $r_{m}$. To this end, we calculated the ground state of BCC phase iron (BCC-Fe) for both ferromagnetic (FM) and anti-ferromagnetic (AFM) configurations using PW and NAO basis sets. The atomic magnetic moments of Fe atom were then estimated using the modulated orbital local projection algorithm with various $r_{m}$. As a reference, we define the “TMAG” by directly summing the magnetic density within the cell $\sum_{\mathbf{r}}\mathbf{m(r)}/N$ for the FM configuration, and “AMAG” as the modulus sum of the magnetic density $\sum_{\mathbf{r}}|\mathbf{m(r)}|/N$ for the AFM configuration, where $N$ is the number of Fe atom. For the NAO calculations, we can use the high-precision TZDP-10 Bohr results as a AFM reference.
Figure 3: Estimation of BCC-Fe atomic magnetic moments as a function of modulation radius $r_{m}$ using the modulated orbital projection algorithm, employing DZP and TZDP basis sets with cutoff radius $r_{c}$ ranging from 6.0 to 10.0 Bohr, to demonstrate the impact of different NAO basis sets on atomic magnetic moments. Where (a) is for the PW basis set in the FM magnetic configuration, with the reference value being the total magnetization per atom (TMAG); (b) is for the NAO basis set in the FM magnetic configuration, with the reference value being the TMAG; (c) is for the PW basis set in the AFM magnetic configuration, with the reference value being the absolute magnetization per atom (AMAG); and (d) is for the LCAO basis set in the AFM magnetic configuration, with the reference value being the atomic magnetization from Mulliken charge.
As shown in Fig. 3(a) and (c), the projected magnetic moments obtained from the different modulated orbitals agree well with each other if $r_{m}<4$ Bohr in all PW calculations. This indicates that the projection results are dependent solely on the selected $r_{m}$, rather than on $r_{c}$. The results from the NAO basis (Fig. 3(b) and (d)) show differences in the projection results between various basis sets, with these discrepancies arising solely from the NAO basis set itself. Furthermore, the comparison between left and right panel of Figs. 3 shows the higher-precision TZDP results align well with the PW ones, while the DZP results ensure qualitative correctness.
The choice of an appropriate modulation radius $r_{m}$ significantly affects the estimated values of atomic magnetic moments. We discuss two possible strategies for determining $r_{m}$. The first approach is to calibrate the projected orbital radius by matching it to reference values, such as experimental data or literature values. However, this can be challenging for complex compounds or systems without available experimental data, as it may be difficult to identify suitable reference values. The second approach is to select the modulation radius corresponding to the maximum atomic magnetic moment. This is based on the rationale that, at this point, the projected orbitals exhibit the largest difference in projection results of wave functions with different spins, thereby providing the most accurate description of atomic magnetic moments and the energy variation caused by spin.
For BCC-Fe, the modulation radius chosen using the first approach is 2.0 Bohr, with the FM and AFM atomic magnetic moments being 2.21 and 1.48 $\mu_{B}$, respectively. These values are in good agreement with those calculated by VASP Ma and Dudarev (2015), which are 2.22 and 1.52 $\mu_{B}$. The second approach determines the radius to be 3.0 Bohr, with the estimated atomic magnetic moments being 2.36 $\mu_{B}$ for FM and 1.6 $\mu_{B}$ for AFM. In this work, we adopted the more general second strategy, selecting 3.0 Bohr as the target radius. Subsequent tests will demonstrate that this choice of projected orbital provides a reliable description of the magnetic potential energy surface and related physical quantities.
III.3 Finite Difference Tests
Figure 4: Finite difference tests for (a) the atomic force in BCC Fe16, (b) the magnetic torque in ternary alloy NiMnTi (c) the cell stress in binary alloy FeNi. Here “Analytic” represents the value calculated from the formula, while “Numerical” is the value determined from the finite difference method. “11” refers to the component of stress matrix $\sigma_{11}$.
Based on the theory defined in Section II, we implemented the spin-cDFT method in the open-source software ABACUS, the method is available with either plane wave or numerical atomic orbital basis sets. In the spin-cDFT framework, we introduced corrections for the energy, atomic forces, and lattice stresses under full magnetic constraints and incorporated calculations of magnetic torque $\bm{\lambda}$. These magnetic torques are optimized in the inner loop of the self-consistent field iterations. To validate our implementation, we compared analytical solutions for atomic forces, lattice stresses, and magnetic torques against numerical ones obtained through finite difference method. As shown in Fig. 4, these tests were conducted on elemental body-centered cubic (BCC) iron (Fe), FeNi binary alloy, and NiMnTi ternary alloy. For a 16-atom BCC-Fe supercell, we perturbed one Fe atom’s position, resulting in a close match between the finite-difference derived forces and the analytical values. In the case of NiMnTi, we perturbed an atom and found high consistency between analytical forces and finite-difference stress tensor values. Crucially, magnetic torques are the partial derivatives of the energy with respect to magnetic moments, rather than atomic positions. Selecting one Fe atom in the FeNi alloy, we perturbed its magnetic moment, which showed perfect agreement between energy-derived magnetic forces and analytically obtained values from the inner loop. For these three tests, the cutoff energy was set to 100 Ry. The Brillouin zone was uniformly sampled at 0.14 Bohr-1 for Fe16 and 0.15 Bohr-1 for FePt/NiMnTi. The successful finite difference tests corroborate the proper functionality of the cDFT features for energy, atomic forces, magnetic torques, and lattice stress in ABACUS.
III.4 BCC and FCC Structures of Iron
Figure 5: Comparison between the unconstrained and constrained DFT results for ferromagnetic BCC-Fe (a) and FCC-Fe (b). The solid lines show the total energy per atom as a function of the cell volume per atom. Here “DZP-8au-v2.1” represents the v2.1 NAOs calculations based on the DZP orbitals with the 8 Bohr cutoff. The gray histograms represent the atomic magnetic moment $m$ after fully unconstrained self-consistent calculations. The dotted lines show the cDFT energy with certain constrains, and plot the corresponding magnetic force $\lambda=|\bm{\lambda}|$ in the upper panel.
The spin-cDFT method presented in this paper facilitates calculations on any magnetic configurations and supports the analysis of complex magnetic structures. We take the bulk iron as an example to demonstrate the reliability of the magnetic constraint method in ABACUS.
The BCC and FCC phases are two prevalent iron crystal structures. The BCC structure is the stable phase of iron at room temperature and is known for its high strength and low ductility. The BCC Fe structure exhibits a ferromagnetic (FM) state with atomic magnetism around 2.2 $\mu_{B}$ per atom Acet et al. (1994), and the Curie temperature was found at 1043 K experimentally Chen and Sundman (2001). On the other hand, The FCC structure has a higher density of atoms compared to BCC. Experimentally, iron undergoes a structural phase transition from BCC to FCC around 1085 K Okamoto et al. (2016) and the FCC structure is the stable phase until 1667 K Okamoto et al. (2016). Although FCC Fe exhibits paramagnetic behavior in the experiment, theoretically, it is predicted that the energy of either antiferromagnetic (AFM) or double-layer antiferromagnetic (DAFM) states would be lower than that of ferromagnetic at 0 K Rinaldi et al. (2024). It has a more unimodal density of states at the Fermi level, which results in lower magnetic moments compared to BCC Fe.
Figure 6: Total energy of BCC-Fe (a) and FCC-Fe (b) as a function of atomic magnetic moment and cell volume. All calculations are performed by spin-cDFT with DZP-8au-v2.1 NAOs orbitals.
Firstly, we consider ferromagnetic BCC-Fe. In the lower panel of Fig. 5(a), the blue solid line illustrates the total energy’s dependence on the BCC-Fe cell volume, calculated using the PW basis without spin constraint. It is evident that the equilibrium volume of BCC-Fe, determined by the volume at which the total energy is minimized, is 11.2 Å3. The black line represents results computed with the v2.1 NAOs basis, specifically the DZP basis with a cutoff radius of 8 au, referred to as “DZP-8au-v2.1”. The NAOs basis is optimized on the basis of the results from the PW basis set, and the energy difference between them reflects the quality of the NAOs orbitals. We observe that the energy difference remains nearly constant under tensile conditions but decreases significantly under compression, indicating discrepancies in orbital precision at varying interatomic distances. The grey histograms in the figure depict the atomic magnetic moments at different volumes. We present only the NAO results since the magnetic moments calculated using the PW basis set and the NAOs basis set are relatively similar. The atomic magnetic moments increase monotonically with volume, with the projected atomic magnetic moment reaching 2.36 $\mu_{B}$ at the energy minimum.
We now impose a constraint on the magnetic configuration, fixing the magnetic moments at 2.36 $\mu_{B}$. The corresponding energy from the spin-cDFT calculations is shown as dotted lines. Due to this constraint, the magnetic moment of 2.36 $\mu_{B}$ represents an excited state for all volume points except 11.2 Å3, resulting in energies higher than the unconstrained ground-state energy. The spin-cDFT energy intersects the ground-state energy at only one point with the magnetic moment 2.36 $\mu_{B}$. The upper panel of Fig. 5(a) displays the magnetic torques $\bm{\lambda}$ optimized in the cDFT calculations. The magnetic torques are zero when the cDFT states are the same as the ground state, indicating that no penalty is required to maintain the specified magnetic moment configuration. However, the excited magnetic states lead to finite magnetic torque, which can be interpreted as an additional effective magnetic field necessary to constrain the magnetic moment.
Similarly, in Fig. 5(b), we present the results for ferromagnetic FCC-Fe. Compared to BCC-Fe, the FCC phase near the ground state exhibits two local minima in the cell volume per atom, located around 10.5 Å3 and 12.2 Å3, respectively. The atomic magnetic moment increases monotonically with the cell volume. A key difference arises at the boundary between the two minima, where a sudden change in the magnitude of the atomic magnetic moment is observed. The atomic magnetic moments corresponding to the two minima are 1.16 $\mu_{B}$ and 2.72 $\mu_{B}$, respectively. We constrained the magnitude of the atomic magnetic moments to these two values, maintaining their direction in the ferromagnetic state. The results show that, regardless of whether the calculations are performed using the PW or NAO basis, the total energy under the constrained conditions only touches the ground-state energy at the corresponding ground-state magnetic moments, with the associated magnetic force $\bm{\lambda}$ being zero. When the constrained magnetic moments deviate from the ground-state values, the total energy exceeds the ground-state energy, and the magnetic force progressively increases.
Figure 7: (a) the BCC-Fe with 2 Fe atoms labels as $\{M_{0}-\Delta M,\theta\}$, where the two atoms have the same atomic magnetic moment $M_{0}-\Delta M$. $M_{0}$ is the atomic magnetic moment and $\Delta M$ is ranging from -0.4 $\mu_{B}$ to 0.4 $\mu_{B}$. $\theta$ is the angle between them. (b-c) shows the magnetic energy barrier difference between the NAOs basis and PW basis, namely $\Delta\hat{E}(M,\theta)=\hat{E}(M,\theta)|_{\rm NAOs}-\hat{E}(M,\theta)|_{\rm PW}$. Here the magnetic energy barrier is defined by the energy above the ground state $\hat{E}(M,\theta)=E(M,\theta)-E(M_{0},0)$. (d-f) shows the ratio of the magnetic energy barrier difference $\Delta\hat{E}(M,\theta)$ to the PW results.
To systematically present the potential energy surfaces of Fe, we modeled BCC-Fe and FCC-Fe with atomic volumes ranging from 8 to 16 Å3 and calculated their total energies at various magnetic moments using spin-cDFT. The magnetic moments considered ranged from 0.2 to 3.8 $\mu_{B}$. Fig. 6 illustrates the potential energy surfaces, $E(V,M)$, for BCC-Fe and FCC-Fe, respectively. The results reveal that the BCC phase exhibits a single energy minimum, whereas the FCC phase features two distinct energy minima, consistent with the findings presented in Fig. 5.
Magnetic calculations have high accuracy requirements. Unlike the PW basis, where accuracy can be systematically improved by increasing the number of plane waves, the accuracy of the NAOs basis depends on the quality of the orbitals themselves and lacks a systematic way to enhance precision. In the following, we will quantitatively examine the precision of spin-cDFT results using different basis sets. In the Heisenberg model $H=\sum_{ij}J_{ij}\mathbf{S}_{i}\cdot\mathbf{S}_{j}$, the magnetic exchange strength $J$ determines the resistance of the material’s magnetic order to external perturbations such as temperature and magnetic fields. This strength typically falls within the range of a few to tens of meV Szilva et al. (2023). The value of $J$ can generally be derived from the energy differences between various magnetic configurations, making the magnetic barrier energy an excellent quantitative metric for evaluating accuracy.
To this end, we chose a BCC-Fe unitcell containing two Fe atoms, each with the same magnetic moment $M$ and an angle $\theta$ between them. When $\theta=0^{\circ}$, the system represents a ferromagnetic state, and when $\theta=180^{\circ}$, it represents an antiferromagnetic state. To evaluate the quality of orbitals, we selected different NAO basis sets with various cutoff radius $r_{c}$ and calculated $E(M,\theta)$ for these configurations. Additionally, we used results from a PW basis as a reference. All $E(M,\theta)$ values for these magnetic configurations were computed using spin-cDFT method. It is important to note that the energy errors calculated using the NAOs basis cannot be directly compared to those of the PW basis. Since the magnetic exchange $J$ reflects the response to changes in magnetic moments, we must first subtract the ground-state energy. Specifically, we define $\hat{E}(M,\theta)=E(M,\theta)-E(M_{0},0)$, where $M_{0}$ is the atomic magnetic moment of the ground state. We then compare the error in $\hat{E}(M,\theta)$ between the NAOs and PW basis.
Fig. 7 presents the results for various NAOs basis sets, showing the energy error compared to the PW results as $\Delta\hat{E}(M,\theta)=\hat{E}(M,\theta)|_{\rm NAOs}-\hat{E}(M,\theta)|_{\rm PW}$. According to the figure, except in regions where the magnetic moment decreases and the angle is relatively large—where the energy error is significantly higher—the error is fairly uniform in other areas. The results indicate that the error primarily depends on the type of basis rather than the cutoff radius. In practical scenarios, such as the gradual transition from a ferromagnetic to a nonmagnetic state with increasing temperature in BCC-Fe, the magnetic moments significantly change in angle instead of amplitude. When the angle changes from $0^{\circ}$ to $180^{\circ}$, the error for the DZP basis set ranges from 2 to 5 meV, whereas the error for the TZDP basis set is significantly reduced to 0.5 to 1 meV. The experimental Curie temperature of BCC-Fe is 1043 K, corresponding to a thermodynamic energy of approximately 90 meV. Given this, both DZP and TZDP basis sets are expected to provide relatively accurate $J$ values, with the precision of the TZDP basis set nearly matching that of the PW basis.
To validate this inference within the framework of density functional theory, we analytically calculated the magnetic exchange energy using the magnetic force theorem Liechtenstein et al. (1984, 1987). Specifically, we utilized the exchange energy calculations by ABACUS and TB2J He et al. (2021), employing a $9\times 9\times 9$ supercell to determine the $J_{ij}$ values up to the fourth-nearest neighbors. The results obtained using different NAOs basis are summarized in Table 1. Since ABACUS currently does not support an interface with PW and TB2J, we used PW results reported in the literature as reference values Wang et al. (2010); Pajda et al. (2001). It is evident that the v2.1 NAOs orbitals are more accurate than those in version 2.0, and the results obtained with the TZDP basis set closely match the reference values, with an error margin within 5%.
Table 1: Exchange energy of ferromagnetic BCC-Fe calculated by various NAOs basis. The cutoff radii is 7 Bohr for all orbitals. “NN” represents the nearest-neighbor exchange, “2NN” refers to the second one, and beyond.
$J_{ij}$ (meV) DZP-v2.0 DZP-v2.1 TZDP-v2.1 Ref. Wang et al.,2010 Ref. Pajda et al.,2001
NN 11.73 17.88 20.23 16.57 19.48
2NN 12.06 11.58 12.23 14.69 11.09
3NN -0.05 -0.36 -1.11 -0.57 -0.20
4NN -2.17 -1.15 -1.93 -2.52 -1.71
III.5 DeePSPIN Model
Figure 8: For (a) BCC and (b) FCC Fe, the comparison of the energy predicted by the DeePSPIN model with those calculated by ABACUS cDFT for the collinear magnetic configurations (FM/AFM/DAFM) with various atomic magnetic moments.
The non-collinear spin-cDFT implementation in ABACUS provides a powerful tool for studying complex magnetic configurations at the atomic scale. However, conventional DFT methods are time consuming, which makes it challenging to simulate magnetic dynamic processes, such as transitions in magnetic ordering that require large-scale supercells. DeePSPIN Yang et al. (2024) is a deep learning approach for magnetic materials that treats spin as so-called “pseudo-atoms” and integrates with the descriptor framework of DeepPot-SE Zhang et al. (2018), preserving translational, rotational, and permutation symmetries. In detail, the DeePSPIN model requires high-precision first-principles data for magnetic materials, including energy, forces, magnetic torques, stress (optional), and atomic configurations, as training data. Additionally, the design of appropriate loss functions and active learning methods Zhang et al. (2020) significantly reduces the required number of samples. A well-trained DeePSPIN model can accurately predict physical properties such as energy, forces, and magnetic torques. By combining it with methods like molecular dynamics (MD) Thompson et al. (2022) and the Landau-Lifshitz-Gilbert equations Gilbert (2004), it can simulate spin evolution in large-scale systems and accurately describe spin-lattice interactions.
Figure 9: For ferromagnetic (a) BCC and (b) FCC Fe, the energy predicted by the DeePSPIN model as a function of volume with various atomic magnetic moments.
To assess the capability of ABACUS+DeePSPIN in representing significant changes in the lattice and magnetic moments, we generated over 10,000 collinear magnetic configurations for the BCC/FCC phase of iron, including three magnetic states of ferromagnetic (FM), antiferromagnetic (AFM), and double-layer ferrimagnetic (DAFM). The cell volume perturbations ranged from -30% to +30%, and the magnitude of the atomic magnetic moments varied from 0 to 4.0 $\mu_{B}$. These data points were obtained through collinear spin-cDFT calculations using the DZP-8au-v2.1 basis set. We then trained a DeePSPIN model based on these perturbed configurations. Fig. 8 adopts both DeePSPIN and DFT methods to predict the total energy of collinear magnetic configurations such as FM, AFM, and DAFM. We find the results from the two models align well with each other, demonstrating the good accuracy of the DeePSPIN model. Fig. 9 illustrates the dependence of total energy for varying magnitudes of magnetic moments, with the volume being changed isotropically. The solid line represents the ground-state energy from unconstrained DFT calculations at each volume, while the data points from the DeePSPIN model correspond to excited states, which have energies higher than those of the DFT results. These observations are in good agreement with recent studies Rinaldi et al. (2024), demonstrating the reliability of the ABACUS spin-cDFT data and the expressive capability of the DeePSPIN model.
Figure 10: Magnetization of BCC-Fe as a function of temperature. Based on the DeePSPIN model, MD simulations at different temperatures are performed, with the stable magnetic moment counted as the magnetization at that temperature. Line width refers to the error bar. The number of atoms in the supercell is 1024, and the virtual magnetic mass M=0.01.
To explore more complex magnetic phenomena, such as the transition of magnetic ordering, it is crucial for the model to accurately capture information about non-collinear magnetic moments. To achieve this, we used ABACUS to generate first-principles data and trained a DeePSPIN model for iron using an active learning strategy Zhang et al. (2020). In detail, we performed random perturbations of the unitcell, atomic positions, and magnetic moments based on the ground-state configurations of BCC-FM/FCC-AFM/DAFM ($2\times 2\times 2$ supercell including 16 Fe atoms). For each configuration, we applied 100 random perturbations, which were used as the initial training data for DPGEN sampling. The maximum perturbation amplitude for the unitcell was 1%, while the maximum perturbation for atomic positions was 0.03 Å. Magnetic moments were randomly rotated by up to 90 degrees, and their magnitudes were perturbed by up to 0.5 $\mu_{B}$. Sampling was conducted separately for BCC and FCC configurations. The sampling temperature ranged from 50 to 1600 K, with each temperature range including MD configurations with the virtual magnetic mass of 0.01/0.05/0.1/0.5/1.0. We employed both NAOs basis (DZP-7au-v2.0) and PW basis to generate two separate data sets. All computations were performed using non-collinear spin-cDFT, incorporating spin-orbit coupling. The PW training strategy followed the same approach as for the NAOs, with the key difference being that NAOs BCC sampling involved 400 configurations per round, whereas PW BCC sampling involved 200 configurations per round. Ultimately, the NAOs basis set generated a total of 34,703 DPGEN samples, while the PW basis set generated 20,931 DPGEN samples. Additionally, for each basis set, we supplemented the data with 175 perturbed configurations for FCC-FM and 109 configurations for the FCC $4\times 2\times 2$ supercells. It is worth noting that DPGEN employs a custom-built version of LAMMPS, designed to be compatible with DeepSpin et al. (2025), for NVT sampling.
We trained two models, DeePSPIN-DZP and DeePSPIN-PW, based on first-principles data, including energy, force, and magnetic torque. These models allow us to perform large-scale simulations to observe the evolution of magnetism with temperature. We conducted NVT ensemble simulations at different temperatures using the trained models, employing an $8\times 8\times 8$ supercell (consisting of 1024 atoms), with a virtual magnetic mass set to 0.01, a timestep of 0.1 fs, for a total of 3 ps of simulation. After the MD equilibration, we selected the trajectory from 1 to 3 ps to calculate the average total magnetization of the system. In Fig.10, we present the total magnetization of BCC-Fe as a function of temperature. It can be observed that BCC-Fe exhibits ferromagnetic behavior at low temperatures, with the total magnetization exceeding 2 $\mu_{B}$. The magnetization decreases gradually up to 800 K, before undergoing a sudden drop around 1000 K, at which point it approaches zero, indicating a ferromagnetic-to-paramagnetic transition. This transition temperature is in close agreement with the experimentally observed Curie temperature of 1043 K. Surprisingly, although the DZP-7au-v2.0 basis set introduces non-negligible errors when calculating the magnetic exchange strength $J$, the magnetization as a function of temperature and the Curie temperature is in very close agreement with the results from PW basis.
In summary, we have implemented the ABACUS spin-cDFT method, which combines first-principles data with AI-assisted magnetic models to extend the precision of magnetic first principles to larger scales, which allows molecular dynamics simulations of lattice and magnetic moment behavior at large spatial scales and long-time scales. We demonstrate that the deep learning magnetic model accurately predicts the Curie temperature of iron. Furthermore, the model can be used to study complex dynamic phenomenon such as the BCC-FCC structural phase transition of iron. A more detailed discussion of this will be provided in a forthcoming paper.
IV Conclusion
In this study, we have implemented a non-collinear spin-constrained method within the open-source software ABACUS, utilizing both plane wave and numerical atomic orbital basis. This implementation allows for the precise control and calculation of arbitrary magnetic configurations. We introduce a smooth modulation orbital method for calculating atomic magnetization using the NAOs projection. Systematic investigations on bulk iron have demonstrated its reliability. The precision of the numerical atomic orbitals used in magnetic calculations has been quantitatively discussed, with the TZDP basis set yielding results very close to the reference plane-wave basis.
The DeePSPIN model trained with the first-principles data generated by this method can predict the energy for different volumes and magnetic moments with high accuracy, demonstrating its good adaptability. Combined with molecular dynamics simulations, the DeePSPIN model successfully observed the ferromagnetic-paramagnetic transition near the experimental Curie temperature.
In conclusion, the noncollinear spin-constrained method implemented in ABACUS provides a powerful tool for studying complex magnetic phenomena at the atomic scale, and a data engine for deep-learning magnetic models. Its integration with AI-assisted magnetic models enables large-scale magnetic simulations, providing new possibilities for the study of complex magnetic phenomena.
Acknowledgements.
We thank Zuxin Jin and Han Wang for many helpful discussions. We gratefully acknowledges AI for Science Institute, Beijing (AISI). M. C. gratefully acknowledges funding support from the National Natural Science Foundation of China (grant no. 12122401,12074007,12135002).